# Universal consistency of the $k$-NN rule in metric spaces and Nagata dimension. III

### v1

@@ 2.4. Implication (3) $\Rightarrow$ (1)
We also note that this implication was established directly in our previous work [4] in the same spirit as the original result of Charles J. Stone for the $k$-NN classifier in the Euclidean space [20].
@@ 3. Nagata dimension
This section is largely (though not entirely) a compilation of known results.
3.1. Nagata sigma-finite dimensional spaces
@@
A metric subspace $X$ of a metric space $\Omega$ has Nagata dimension $\leq\delta\in{\mathbb{N}}$ on the scale $s>0$ inside of $\Omega$ if every finite family of closed balls in $\Omega$ with centres in $X$ admits a subfamily of multiplicity $\leq\delta+1$ in $\Omega$ which covers all the centres of the original balls. The subspace $X$ has a finite Nagata dimension in $\Omega$ if $X$ has finite dimension in $\Omega$ on some scale $s>0$. Notation for the smallest $\delta$ with the above property: $\dim^{s}_{Nag}(X,\Omega)$ or sometimes simply $\dim_{Nag}(X,\Omega)$.
Following Preiss [18], let us call a family of balls disconnected if the centre of each ball does not belong to any other ball. Here is a mere reformulation of the above definition.
@@
In the definition 3.1, as well as in the proposition 3.1, it is convenient to replace closed balls with open ones.
@@
Let $\gamma$ be a finite disconnected family of open balls in $\Omega$ of radii <s<s, centred in $\bar{X}$. Let $y\in\Omega$, and let $\gamma^{\prime}$ consist of all balls in $\gamma$ containing $y$. Choose ${\varepsilon}>0$ so small that the open $\varepsilon$}-ball around $y$ is contained in every element of $\gamma^{\prime}$. For every open ball $B\in\gamma^{\prime}$, denote $y_{B}$ the centre and $r_{B}$ the radius. We can additionally assume that ${\varepsilon}<r_{B}$. Denote $B^{\prime}$ an open ball of radius $r_{B}-{\varepsilon}>0$, centred at a point $x_{B}\in X$ satisfying $d(x_{B},y_{B})<{\varepsilon}/2$. Note that $y\in B^{\prime}$ by the triangle inequality. Again by the triangle inequality, the family $\{B^{\prime}\colon B\in\gamma^{\prime}\}$ is disconnected, and it has radii <s<s. Therefore, $y$ only belongs to $\leq\delta+1$ balls $B^{\prime}$, $B\in\gamma^{\prime}$. Consequently, the cardinality of $\gamma^{\prime}$ is bounded by $\delta+1$. ∎
@@
Due to Proposition 3.1, in the above definition we can assume the subspaces $X_{n}$ to be closed in $\Omega$. By Proposition 3.1, we can assume $X_{n}$ to form an increasing chain.
@@
Let us say, slightly extending the definition 3.1 of Preiss, that a subset $A$ of $\Omega$ is sigma-finite dimensional in $\Omega$ if one can represent $A$ as the union of a countable family of subsets $A_{n}$ each of which is finite-dimensional in $\Omega$ on some positive scale.
Denote $\gamma$ the family of all non-empty open subsets of $\Omega$ that are sigma-finite dimensional in $\Omega$. Let $O$ be the union of $\gamma$. If it happens that $O=\emptyset$, there is nothing left to prove, so we may as well assume $\gamma\neq\emptyset$. Since $O$ is a separable metric space, it is Lindelöf, and one can select a countable subfamily of $\gamma$ whose union is $O$ ([8], Thm. 3.8.1). It follows that $O$ itself is sigma-finite dimensional in $\Omega$, so it is the largest open subset sigma-finite dimensional in $\Omega$.
@@
In [18], Preiss seems to imply (p. 567, lines 3-4) that every complete separable metric space $\Omega$ that is not sigma-finite dimensional (in itself) in the sense of Nagata contains a non-empty closed subspace $\Upsilon$ containing no non-empty (relatively) open subsets of finite Nagata dimension in $\Upsilon$ proper, but I am unable to verify this claim. The purported proof of it that my student Sushma Kumari and I had come up with and included in her Ph.D. thesis (Lemma 1.4.2, p. 35 in [10]) appears deficient.
3.2. Further remarks
{rmrk}
A metric space is said to have finite Nagata dimension if it has finite Nagata dimension in itself. The Nagata–Ostrand theorem [13, 14, 16] states that the Lebesgue covering dimension of a metrizable topological space is the smallest Nagata dimension of a compatible metric on the space (and in fact this is true on every scale $s>0$). This is the historical origin of the concept of the Nagata metric dimension.
@@ 3.2. Further remarks
It is easy to see, especially using Prop. 3.1, that $\dim_{Nag}({\mathbb{R}})=1$.
@@
In ${\mathbb{R}}^{2}={\mathbb{C}}$ the family of closed balls of radius one centred at the vectors $\exp(2\pi ki/5)$, $k=1,2,\ldots,5$, has multiplicity $5$ and admits no proper subfamily containing all the centres. Therefore, the Nagata dimension of $\ell^{2}(2)$ is at least $5$. Since the plane can be covered with $6$ cones having the central angle $\pi/3$, Example 3.2 implies that $\dim_{Nag}(\ell^{2}(2))=5$.
@@
4. The measure construction
4.1. The indexing tree
Let $\Omega$ be a separable complete metric space that is not sigma-finite dimensional in the sense of Nagata. Fix a closed subspace $\Upsilon\subseteq\Omega$ with the property that none of its non-empty relatively open subsets has finite Nagata dimension in $\Omega$ (Proposition 3.1).
We will work with an unbounded sequence $(m_{i})_{i=0}^{\infty}$ of positive natural numbers starting with $m_{0}=1$, which is arbitrary at this stage, but will be recursively specified at a later point. The only condition we impose from the start is that $m_{i}\geq 2$ for $i\geq 1$. Form a rooted prefix tree, $T$, with vertices being finite words $t=t_{1}t_{2}\ldots t_{i}$, $i\in{\mathbb{N}}$, $t_{j}\in[m_{j}]=\{1,2,\ldots,m_{j}\}$, $j=1,2,\ldots,i$. For such a $t$, we denote $i=\lvert t\rvert$. The root is the empty word, denoted $\ast$; so, $\lvert\ast\rvert=0$. In this way, the children of the node $t$ with $\lvert t\rvert=i$ are the words $t1,t2,\ldots,tm_{i+1}$.
4.2. Selecting points $x^{t},y^{t}$
@@ 4.2.1. Base of recursion
At the first step, we use the fact that $\Upsilon$ itself is not Nagata finite dimensional in $\Omega$. In particular, as $\dim^{s}_{Nag}(\Upsilon,\Omega)>m_{1}$ on any scale $s>0$, there exists a disconnected collection of $m_{1}$ balls of radii <s<s centred in $\Upsilon$ (which balls we can take open by Proposition 3.1), having a point in $\Omega$ in common. At the base step, the choice of a scale does not matter, we can take $s=1$.
Denote the centres of the selected balls $y^{1},y^{2},\ldots,y^{m_{1}}\in\Upsilon$, the radii $r_{1}^{1},\ldots,r_{1}^{m_{1}}>0$, and the common point of the open balls $x^{\ast}\in\Omega$. Take ${\epsilon}^{\ast}>0$ so small that ${\epsilon}^{\ast}<1$ and for each $1,2,\ldots,m_{1}, iji\neq j$, we have
@@ 4.2.2. Step of recursion
Now, let $i\geq 1$ and suppose that $x^{t}\in\Omega$ and ${\epsilon}^{t}>0$ have been selected for all $t$ with $0\leq\lvert t\rvert<i$, and $y^{t}\in\Upsilon$ have been chosen for all $t$ with $1\leq\lvert t\rvert\leq i$, in a way to satisfy the properties (1)-(3) in subs. 4.2.
@@
4.3. Compact set $K$ supporting two measures
By the condition 4.2(3), for any $t$ with $\lvert t\rvert=i\geq 1$ and any word $s$, we get $d(y^{t},y^{ts})<2^{-i+1}$. Also, the condition 4.2(2) says that $d(y^{t},x^{t})<2^{-i+1}$. Consequently, all the points $y^{t}$, $x^{t}$ with $\lvert t\rvert\geq i$ are contained within the distance $2^{-i+2}$ from the finite set of points $y^{t}$, $\lvert t\rvert=i$. This implies the entire countable set of points $x^{t}$, $\lvert t\rvert\geq 0$, $y^{t}$, $\lvert t\rvert\geq 1$ is precompact. Denote $K$ its closure, a compact subset of $\Omega$.
We will now construct two finite measures supported on $K$, each one of total mass $1/2$: a purely atomic measure $\mu_{1}$, having $x^{t}$, $t\in T$ as atoms, and a non-atomic measure, $\mu_{0}$, which is a weak limit of the sequence of uniform measures supported on the points $y^{t}$, $\lvert t\rvert=i$, as $i\to\infty$.
@@ 4.3.2. Measure $\mu_{0}$
The sequence of measures $(\mu_{0}^{(i)})$ converges in the weak topology in the space of finite measures on the compact space $K$. Indeed, the weak topology (that is, the weak topology on the dual space of $C(K)$) on the measures of fixed mass on a compact space is metrized by the mass transportation distance (see e.g. [21], corollary 6.13). For every $t$, $\lvert t\rvert\geq 1$, the $\mu_{0}^{(i)}$-measure of the ball $\bar{B}_{{\epsilon}^{t^{-}}}(y^{t})$ is equal to $1/(2\prod_{j\leq i}m_{i})$, because the ball only contains exactly one point among all $y^{t}$ with $\lvert t\rvert=i$, namely, its centre. If now $j>i$, then the number of points $y^{s}$ with $\lvert s\rvert=j$ contained in the same ball $\bar{B}_{{\epsilon}^{t^{-}}}(y^{t})$ is equal to $\prod_{\ell=i+1}^{j}m_{\ell}$, and so the total mass of the ball under the measure $\mu_{0}^{(j)}$ is the same as under $\mu_{0}^{(i)}$. The transport of mass between the restrictions of the two measures on the ball requires shifting one of them by a distance $\leq{\epsilon}^{t^{-}}<2^{-i}$. We conclude that the overall mass transportation distance between $\mu_{0}^{(i)}$ and $\mu_{0}^{(j)}$ is bounded by $2^{-i}$ times the total mass, $1/2$. This way, the sequence $(\mu_{0}^{(i)})$ is Cauchy, hence convergent in the weak topology on the compact space of measures of mass $1/2$. Denote the limit $\mu_{0}$. It also satisfies $\mu_{0}(K)=1/2$.
For every $i$, the union of the closed balls $\bar{B}_{{\epsilon}^{t^{-}}}(y^{t})$, $\lvert t\rvert=i$ has full $\mu_{0}$-measure, $1/2$. In particular, it follows that the measure $\mu_{0}$ is diffuse, because for each $t$ with $\lvert t\rvert=i$, it has the property $\mu_{0}(\bar{B}_{{\epsilon}^{t^{-}}}(y^{t}))=1/(2\prod_{0\leq j\leq i}m_{i})\to 0$ as $i\to\infty$.
@@ 4.3.3. The probability measure $\mu$ and the regression function $\eta$
This $(\mu,\eta)$ will become our learning problem. The Bayes (optimal) classifier for this problem is a.e. equal to the binary function $\eta$, so will assume the value $1$ with probability half. As we will see, under a suitable selection of the sequence $(m_{i})$ used to construct $\mu$ and $\eta$, the $k$-NN classifier with a given sequence of values $(k_{n})_{n=1}^{\infty}$ does not asymptotically converge to $\eta$ and so is not consistent under this problem.
5. Inconsistency of the $k$-NN under $(\mu,\eta)$
5.1. Selecting $(m_{i})$ and $(n_{i})$
Let the sequence $(k_{n})_{n=1}^{\infty}$ of values of $k$ be given, satisfying $k_{n}\to\infty$ and $k_{n}/n\to 0$. Fix a summable sequence of values of the risk $\delta_{i}>0,i\in{\mathbb{N}}$. Remember that $m_{0}=1$. This is the base of the recursion.

### v2

@@ 2.4. Implication (3) $\Rightarrow$ (1)
We also note that this implication was established directly in our previous work [8] in the same spirit as the original result of Charles J. Stone for the $k$-NN classifier in the Euclidean space [30], though with some interesting geometrical complications.
@@ 3. Nagata dimension
This Section is largely (though not entirely) a compilation of known results. Since they are scattered in the literature [3, 8, 21, 23, 24, 25, 26, 28] and for some of them the proofs are missing, we are giving a self-contained presentation.
@@
A metric subspace $X$ of a metric space $\Omega$ has Nagata dimension $\leq\delta\in{\mathbb{N}}$ on the scale $s\in{\mathbb{R}}\cup\{+\infty\}$ inside of $\Omega$ if every finite family of closed balls in $\Omega$ with centres in $X$ and radii <s<s admits a subfamily of multiplicity $\leq\delta+1$ in $\Omega$ which covers all the centres of the original balls. The subspace $X$ has a finite Nagata dimension in $\Omega$ if $X$ has finite dimension in $\Omega$ on some scale $s>0$. Notation for the smallest $\delta$ with the above property: $\dim^{s}_{Nag}(X,\Omega)$ or sometimes simply $\dim_{Nag}(X,\Omega)$.
Following Preiss [28], let us call a family of balls disconnected if the centre of each ball does not belong to any other ball from the family. Here is a reformulation of the above definition.
@@
In the definition 3, as well as in the proposition 3, it is convenient to replace closed balls with open ones.
@@
Let $\gamma$ be a finite disconnected family of open balls in $\Omega$ of radii <s<s, centred in $\bar{X}$. Let $y\in\Omega$, and let $\gamma^{\prime}$ consist of all balls in $\gamma$ containing $y$. Choose ${\varepsilon}>0$ so small that the open $\varepsilon$}-ball around $y$ is contained, as a subset, in every element of $\gamma^{\prime}$. For every open ball $B\in\gamma^{\prime}$, denote $y_{B}$ the centre and $r_{B}$ the radius. We can additionally assume that ${\varepsilon}<r_{B}$. Denote $B^{\prime}$ an open ball of radius $r_{B}-{\varepsilon}>0$, centred at a point $x_{B}\in X$ satisfying $d(x_{B},y_{B})<{\varepsilon}/2$. Note that $y\in B^{\prime}$ by the triangle inequality. Again by the triangle inequality, the family $\{B^{\prime}\colon B\in\gamma^{\prime}\}$ is disconnected, and it has radii <s<s. Therefore, $y$ only belongs to $\leq\delta+1$ balls $B^{\prime}$, $B\in\gamma^{\prime}$. Consequently, the cardinality of $\gamma^{\prime}$ is bounded by $\delta+1$. ∎
@@
Due to Proposition 3, in the above definition we can assume the subspaces $X_{n}$ to be closed in $\Omega$. By Proposition 3, we can assume $X_{n}$ to form an increasing chain.
{rmrk}
The Baire category theorem implies that every sigma-finite dimensional metric space $\Omega$ contains a non-empty open subset $X$ having a finite metric dimension in $\Omega$.
{rmrk}
Note that the notion of a subset $X$ of a metric space $\Omega$ being finite-dimensional as studied above is relative, that is, it is a property of the pair $(X,\Omega)$ rather than of $X$ itself.
{rmrk}
A metric space $\Omega$ is said to have finite Nagata dimension if it has finite Nagata dimension in itself, that is, as a pair $(\Omega,\Omega)$. In this case we write $\dim_{Nag}(\Omega)$. The Nagata–Ostrand theorem [23, 24, 26] states that the Lebesgue covering dimension of a metrizable topological space is the smallest Nagata dimension of a compatible metric on the space (and in fact this is true on every scale $s>0$). This is the historical origin of the concept of the Nagata metric dimension.
@@
Let us say, slightly extending the definition 3 of Preiss, that a subset $A$ of $\Omega$ is sigma-finite dimensional in $\Omega$ if one can represent $A$ as the union of a countable family of subsets $A_{n}$ each of which is finite-dimensional in $\Omega$ on some strictly positive scale.
Denote $\gamma$ the family of all non-empty open subsets of $\Omega$ that are sigma-finite dimensional in $\Omega$. Let $O$ be the union of $\gamma$. If it happens that $O=\emptyset$, there is nothing left to prove, so we may as well assume $O\neq\emptyset$. Since $O$ is a separable metric space, it is Lindelöf, and one can select a countable subfamily of $\gamma$ whose union is $O$ ([14], Thm. 3.8.1). It follows that $O$ itself is sigma-finite dimensional in $\Omega$, so it is the largest open subset that is sigma-finite dimensional in $\Omega$.
@@
In [28], Preiss seems to imply (p. 567, lines 3-4) that every complete separable metric space $\Omega$ that is not sigma-finite dimensional (in itself) in the sense of Nagata contains a non-empty closed subspace $\Upsilon$ containing no non-empty (relatively) open subsets of finite Nagata dimension in $\Upsilon$ proper, but I am unable to verify this claim. The purported proof of it that my student Sushma Kumari and I had come up with and she included in her Ph.D. thesis (Lemma 1.4.2, p. 35 in [20]) appears deficient.
@@
It is easy to see, especially using Prop. 3, that $\dim_{Nag}({\mathbb{R}})=1$, for the real line $\mathbb{R}$} with the standard metric.
@@
In ${\mathbb{R}}^{2}={\mathbb{C}}$ the family of closed balls of radius one centred at the vectors $\exp(2\pi ki/5)$, $k=1,2,\ldots,5$, has multiplicity $5$ and admits no proper subfamily containing all the centres. Therefore, the Nagata dimension of $\ell^{2}(2)$ is at least $5$. Since the plane can be covered with $6$ cones having the central angle $\pi/3$, Example 3 implies that $\dim_{Nag}(\ell^{2}(2))=5$.
@@
4. Proof of Theorem 1
4.1. The construction of the learning problem $(\mu,\eta)$
4.1.1. The indexing tree
Let $\Omega$ be a separable complete metric space that is not sigma-finite dimensional in the sense of Nagata. Fix a non-empty closed subspace $\Upsilon\subseteq\Omega$ with the property that none of its non-empty relatively open subsets has finite Nagata dimension in $\Omega$ (Proposition 3).
We will work with an unbounded sequence $(m_{i})_{i=0}^{\infty}$ of positive natural numbers starting with $m_{0}=1$, which is arbitrary at this stage, but will be recursively specified at a later point. The only condition we impose from the start is that $m_{i}\geq 2$ for $i\geq 1$. Form a rooted prefix tree, $T$, with vertices being finite words $t=t_{1}t_{2}\ldots t_{i}$, $i\in{\mathbb{N}}$, $t_{j}\in[m_{j}]=\{1,2,\ldots,m_{j}\}$, $j=1,2,\ldots,i$. For such a $t$, we denote $i=\lvert t\rvert$. The root is the empty word, denoted $\ast$; so, $\lvert\ast\rvert=0$. In this way, the children of the node $t$ with $\lvert t\rvert=i$ are the words $t1,t2,\ldots,tm_{i+1}$. (Here for example $t1$ means the word obtained by concatenating the word $t$ with the letter $1$, etc.)
4.1.2. Selecting points $x^{t},y^{t}$
@@ 4.1.3. Base of recursion
At the first step, we use the fact that $\Upsilon$ itself is not Nagata finite dimensional in $\Omega$. In particular, as $\dim^{s}_{Nag}(\Upsilon,\Omega)>m_{1}$ on each scale $s>0$, there exists a disconnected collection of $m_{1}$ balls of radii <s<s centred in $\Upsilon$ (which balls we can take open by Proposition 3), having a point in $\Omega$ in common. At the base step, we take $s=1$.
Denote the centres of the selected balls $y^{1},y^{2},\ldots,y^{m_{1}}\in\Upsilon$, the radii $0<r_{1}^{1},\ldots,r_{1}^{m_{1}}<1$, and the common point of the open balls $x^{\ast}\in\Omega$. Take ${\epsilon}^{\ast}>0$ so small that ${\epsilon}^{\ast}<1$ and for each $1,2,\ldots,m_{1}, iji\neq j$, we have
@@ 4.1.4. Step of recursion
Now, let $i\geq 1$ and suppose that $x^{t}\in\Omega$ and ${\epsilon}^{t}>0$ have been selected for all $t$ with $0\leq\lvert t\rvert<i$, and $y^{t}\in\Upsilon$ have been chosen for all $t$ with $1\leq\lvert t\rvert\leq i$, in a way to satisfy the properties (1)-(3) in subs. 4.1.2.
@@
4.1.5. Compact set $K$
By the condition 4.1.2(3), for any $t$ with $\lvert t\rvert=i\geq 1$ and any word $s$, we get $d(y^{t},y^{ts})<2^{-i+1}$. Also, the condition 4.1.2(2) says that $d(y^{t},x^{t})<2^{-i+1}$. Consequently, all the points $y^{t}$, $x^{t}$ with $\lvert t\rvert\geq i$ are contained within the distance $2^{-i+2}$ from the finite set of points $y^{t}$, $\lvert t\rvert=i$. This implies the entire countable set of points $x^{t}$, $\lvert t\rvert\geq 0$, $y^{t}$, $\lvert t\rvert\geq 1$ is precompact. Denote the closure of this set $K$. It is a compact subset of the complete metric space $\Omega$.
We will now construct two finite measures both having $K$ as their support, each one of total mass $1/2$. The measure $\mu_{1}$ is purely atomic and has $x^{t}$, $t\in T$ as atoms. The non-atomic measure $\mu_{0}$ is a weak limit of the sequence of uniform measures supported on the points $y^{t}$, $\lvert t\rvert=i$, as $i\to\infty$.
@@ 4.1.7. Measure $\mu_{0}$
The sequence of measures $(\mu_{0}^{(i)})_{i=1}^{\infty}$ converges in the weak topology in the space of finite measures on the compact space $K$. Indeed, the weak topology (that is, the weak topology on the dual space of $C(K)$) on the measures of fixed mass on a compact space is metrized by the mass transportation distance (see e.g. [33], corollary 6.13). For every $t$, $\lvert t\rvert\geq 1$, the $\mu_{0}^{(i)}$-measure of the ball $\bar{B}_{{\epsilon}^{t^{-}}}(y^{t})$ is equal to $1/(2\prod_{j\leq i}m_{i})$, because the ball only contains exactly one point among all $y^{t}$ with $\lvert t\rvert=i$, namely, its centre. If now $j>i$, then the number of points $y^{s}$ with $\lvert s\rvert=j$ contained in the same ball $\bar{B}_{{\epsilon}^{t^{-}}}(y^{t})$ is equal to $\prod_{\ell=i+1}^{j}m_{\ell}$, and so the total mass of the ball under the measure $\mu_{0}^{(j)}$ is the same as under $\mu_{0}^{(i)}$. The transport of mass between the restrictions of the two measures on the ball requires shifting one of them by a distance $\leq{\epsilon}^{t^{-}}<2^{-i}$. We conclude that the overall mass transportation distance between $\mu_{0}^{(i)}$ and $\mu_{0}^{(j)}$ is bounded by $2^{-i}$ times the total mass, $1/2$. This way, the sequence $(\mu_{0}^{(i)})$ is Cauchy, hence convergent in the weak topology on the compact space of measures of mass $1/2$. Denote the limit $\mu_{0}$. It also satisfies $\mu_{0}(K)=1/2$.
For every $i$, the union of the closed balls $\bar{B}_{{\epsilon}^{t^{-}}}(y^{t})$, $\lvert t\rvert=i$ has full $\mu_{0}$-measure, $1/2$. In particular, it follows that the measure $\mu_{0}$ is diffuse, because for each $t$ with $\lvert t\rvert=i$, it has the property $\mu_{0}(\bar{B}_{{\epsilon}^{t^{-}}}(y^{t}))=1/(2\prod_{0\leq j\leq i}m_{i})\to 0$ as $i\to\infty$. (Remembering that $m_{i}\geq 2$.)
@@ 4.1.8. The probability measure $\mu$ and the regression function $\eta$
This $(\mu,\eta)$ will become our learning problem. The Bayes (optimal) classifier for this problem is a.e. equal to the binary function $\eta$, so assumes the values 0 and $1$ with probability half each. As we will see, given any sequence of values $(k_{n})_{n=1}^{\infty}$ of values of $k$ with $k_{n}/n\to 0$, under a suitable selection of the sequence $(m_{i})$ used to construct $\mu$ and $\eta$, the $k$-NN rule with $k=k_{n}$ does not asymptotically converge to $\eta$ in probability.
Namely, for a certain subsequence $(n_{i})$ of the values of $n$, the sequence of the predictions of the $k$-NN classifier under $(\mu,\eta)$, corresponding to the values $n=n_{i}$ and $k=k_{n_{i}}$, will converge in probability, as $i\to\infty$, to the constant predictor $1$. This, the $k$-NN classifier will not be consistent under this problem.
4.2. Choosing the sequences $(m_{i})$ and $(n_{i})$
Let a sequence $(k_{n})_{n=1}^{\infty}$ of values of $k$ be given, satisfying $k_{n}/n\to 0$. Now we will select the sequences $(m_{i})$ and $(n_{i})$.
Fix a summable sequence of values of the risk $\delta_{i}>0,i\in{\mathbb{N}}$. Remember that $m_{0}=1$. This is the base of the recursion.